# Optimal e-value testing for properly constrained hypotheses

### v1

@@ 6 Domains with finite cardinality
Lemma 8 and Lemma 3 ensure that the optimal e-class exists. Hence, by Lemma 2, it is sufficient to show that $\mathcal{E}_{\mathcal{H}}^{\vee}$ coincides with the set of all maximal e-variables. By Lemma 10, we know that all the elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal. Therefore, we are left to show that every maximal e-variable is in $\mathcal{E}_{\mathcal{H}}^{\vee}$. Fix a maximal e-variable $E$. We will adopt the following proof strategy. First, we show that there is a fully supported $\tilde{P}\in\mathcal{H}$ such that $\langle\tilde{P},E\rangle=1$. Using that such $\tilde{P}$ is in the relative interior of $\mathcal{H}$, we will establish that $\langle P,E\rangle=1$ for all $P\in\mathcal{H}$. Finally, we will see that this property is enough to ensure that $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$.
@@
Since this is the maximum among a finite number of continuous functions, it is continuous. For $\varepsilon$ large enough, $\varphi_{x}(\varepsilon)>1$. Since $\varphi_{x}(0)<1$, there is $\tilde{\varepsilon}>0$ such that $\varphi_{x}(\tilde{\varepsilon})=1$. Let $\tilde{E}=E+\tilde{\varepsilon}\mathbf{1}_{x}$. Clearly, for any $\hat{P}\in\mathcal{V}_{0}$ we have $\langle\hat{P},\tilde{E}\rangle=\langle\hat{P},E\rangle=1$, since $x\notin\operatorname*{Supp}(\hat{P})$. Moreover, for all $\hat{P}\in\mathcal{V}\setminus\mathcal{V}_{0}$, $\langle\hat{P},\tilde{E}\rangle\leq 1$, by definition of $\varphi_{x}$ and $\tilde{\varepsilon}$. Therefore, $\tilde{E}$ is an e-variable, since its expectation is at most one under all the vertices of $\mathcal{H}$. As $\tilde{\varepsilon}>0$, $\tilde{E}\succ E$, and so $E$ cannot be maximal, which is a contradiction. We conclude that $S=\mathcal{X}$.
@@ 7 Matching sets
By definition of compatible set, $\mathcal{H}_{S}$ is non-empty, and hence a hypothesis on $S$. We list here a few results outlining basic properties of the restrictions of finitely and properly constrained hypotheses. The proofs, rather short, can be found in Appendix A.
@@ 8 Optimality of the dual e-class
As all elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal (Lemma 10), by Lemma 2 it is enough to show that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. Fix $E\in\mathcal{E}_{\mathcal{H}}$. By Lemma 5 we can fix a minimal constraint $\Phi$ for $\mathcal{H}$. By Lemma 17, there is a finite matching set $S^{\star}\subseteq\mathcal{X}$ such that any closed set $S\subseteq\mathcal{X}$ containing $S^{\star}$ is a matching set that induces a compact $\Lambda_{\Phi|S}$. Since $\mathcal{X}$ is closed, it is sigma-compact, and we can find a sequence $(K_{i})_{i\geq 1}$ of compact matching sets, such that $S^{\star}\subseteq K_{i}\subseteq K_{i+1}$, for all $i\geq 1$, and $\mathcal{X}=\bigcup_{i=1}^{\infty}K_{i}$. By construction, $\mathcal{H}_{K_{i}}$ is properly constrained (Lemma 13) and $\Phi\big{|}_{K_{i}}$ a constraint for it (Lemma 12). From Proposition 4, for each $i\geq 1$ there is $\lambda_{i}\in\Lambda_{\Phi|K_{i}}$ such that $E_{\lambda_{i}}(x)\geq E(x)$ for all $x\in K_{i}$, where as usual $E_{\lambda}=1-\lambda\cdot\Phi$. For all $i\geq 1$, because $K_{i}\subseteq K_{i+1}$ we have $\Lambda_{\Phi|K_{i}}\supseteq\Lambda_{\Phi|K_{i+1}}$. In particular, for any $j\geq 1$, $(\lambda_{i})_{i\geq j}$ is contained in $\Lambda_{\Phi|K_{j}}$, which is compact by Lemma 17. So, there is a subsequence $(\lambda^{\prime}_{i})_{i\geq 1}$ of $(\lambda_{i})_{i\geq 1}$ that converges to $\lambda^{\star}\in\bigcap_{j\geq 1}\Lambda_{\Phi|K_{j}}$. Let $(K^{\prime}_{i})_{i\geq 1}$ be the corresponding subsequence of $(K_{i})_{i\geq 1}$, obtained selecting the same indices. By Lemma 15, $\lambda^{\star}\in\Lambda_{\Phi}$, so $E_{\lambda^{\star}}\in\mathcal{E}_{\mathcal{H}}^{\vee}$. Fix any $x\in\mathcal{X}$. There is an index $i^{\star}\geq 1$ such that $x\in K_{i}^{\prime}$ for all $i\geq i^{\star}$. Thus, $E_{\lambda^{\prime}_{i}}(x)\geq E(x)$ for all $i\geq i^{\star}$, and so $E_{\lambda^{\star}}(x)\geq E(x)$. As the choice of $x$ has been arbitrary, $E_{\lambda^{\star}}\succeq E$. We conclude that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. ∎
@@ 9.2 Loosely constrained hypotheses
For any properly loosely constrained hypothesis the optimal e-class exists. This is made explicit in the next proposition (see Appendix A for the proof).
@@
We remark that the set =$\mathcal{H}=\{P\in\mathcal{P}_{\Phi}\,:\,\langle P,\Phi\rangle=0\}$, where $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ is a continuous function, is non-empty (and thus a hypothesis) if, and only if, $0\in\operatorname*{conv}\Phi(\mathcal{X})$ (by Lemma 22 and Lemma 24). Conversely, when we allow for inequality constraints, as in the definition of loosely constrained hypotheses, this is not anymore a necessary condition. More explicitly, =$\mathcal{H}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\,:\,\langle P,\Phi^{\prime}\rangle=0\,,\,\langle P,\Phi^{\prime\prime}\rangle\geq 0\}$ can be non empty if $0\notin\operatorname*{conv}\Phi(\mathcal{X})$, (where $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ are continuous functions on $\mathcal{X}$ and Φ=($\Phi=(\Phi^{\prime},\Phi^{\prime\prime})$). However, here we only focus on the case $0\in\operatorname*{conv}\Phi(\mathcal{X})$, as required by Definition 10, and defer to future work the analysis of the setting of generic hypotheses defined via loose constraints. We also note that Lemma 19 cannot be trivially extended to the case of loosely constrained hypotheses, and it is not always possible to restrict the domain $\mathcal{X}$ of a loosely constrained hypothesis to make it properly loosely constrained. For an example where this is not possible, consider the case of a loosely constrained hypothesis on $\mathcal{X}=\mathbb{R}$, with tight constraint $\Phi^{\prime}:x\mapsto 1-x$ and slack constraint $\Phi^{\prime\prime}:x\mapsto x^{2}-1$.
@@ 10.1 Mean estimation for bounded random variables
Clearly, $\mathcal{H}_{\mu}$ is a hypothesis on $\mathcal{X}$ (it is non-empty by Lemma 22) and it is finitely constrained, with $\Phi_{\mu}:x\mapsto\mu-x$ a constraint (as $0\in\operatorname*{conv}\Phi_{\mu}(\mathcal{X})$ because $\mu\in\mathcal{C}$). Moreover, if $\mu\in\operatorname*{rel\,int}\mathcal{C}$, $\mathcal{H}_{\mu}$ is a properly constrained hypothesis on $\mathcal{X}$. For this reason, for the sake of simplicity here we assume as known that $\mu^{\star}\in\operatorname*{rel\,int}\mathcal{C}$. Therefore, we are only interested in confidence sets $S_{t}\subseteq\operatorname*{rel\,int}\mathcal{C}$. We note, however, that the general case can be addressed by dealing with the values of $\mu$ on the relative boundary of $\mathcal{C}$ as discussed in Section 9.1. A valid confidence sequence for $\mu^{\star}\in\operatorname*{rel\,int}\mathcal{C}$ with confidence level $\delta\in(0,1)$ is $(S_{t})_{t\geq 1}$, where
@@
Here, $R_{t}(\mu)$ denotes the cumulative reward, up to round $t$, of a player who is playing a testing game as in Definition 2 on the null hypothesis $P^{\star}\in\mathcal{H}_{\mu}$, observing the sequence $(X_{i})_{i\geq 1}$. From the fact that $\{\mu^{\star}=S_{t}\,,\,\forall t\geq 1\}\iff\{R_{t}(\mu^{\star})\leq\log(1/\delta)\,,\,\forall t\geq 1\}$, which happens with probability at least $1-\delta$ by Lemma 1, it follows that $(S_{t})_{t\geq 1}$ is a confidence sequence.
@@ 10.2 Mean estimation for heavy-tailed distributions
where we note that the fact that $\langle P,X^{2}\rangle\leq 1$ automatically implies that $\langle P,X\rangle$ is well defined. We remark that $\mathcal{H}_{1}=\{\delta_{1}\}$ and $\mathcal{H}_{-1}=\{\delta_{-1}\}$. On the other hand, it is easy to check that, for $\mu\in(-1,1)$, $\mathcal{H}_{\mu}$ is a loosely properly constrained hypothesis on $\mathbb{R}$, where $\Phi^{\prime}_{\mu}:x\mapsto\mu-x$ is a tight constraint and $\Phi^{\prime\prime}:x\mapsto 1-x^{2}$ is a slack constraint. Denoting as $R_{t}(\mu)$ the cumulative reward (up to round $t$) of a player betting in a testing game on $\mathcal{H}_{\mu}$ and observing the sequence $(X_{i})_{i\geq 1}$, with the same reasoning as in Section 10.1, we deduce that
@@
Once more, we shall restrict the testing games to the optimal e-classes. By Proposition 5, for $\mu\in(-1,1)$ the optimal e-class is
@@
We note that the problem of finding confidence sequences for heavy-tailed random variables via betting strategies has been addressed in the literature. Agrawal et al. (2021) obtained anytime-valid lower bounds for real random variables with bounded $(1+\varepsilon)$-th (non-centred) moment, via an approach that implicitly uses e-variables, as their Lemma 17 is in fact a regret upper bound for the cumulative reward of a player in a testing game as in Definition 2. Notably, our results show that their analysis considered testing games restricted to the optimal e-class. On the other hand, Wang and Ramdas (2023) leveraged ideas from Catoni (2012) to set up testing games in order to find confidence sequences for the mean of real random variables with bounded $(1+\varepsilon)$-th central moment. However, the theory that we have developed in this work shows that their choice of restriction for the e-variables does not correspond to the optimal e-class.
@@ 11 Conclusion
First, it is worth noticing that key concepts of maximal e-variable, and majorising and optimal e-class that we discuss in this work are strongly connected to the classical statistical notions of admissibility and completeness (see, e.g., Section 1.8 in Lehmann and Romano, 2022). Indeed, with this terminology, a maximal e-variable can be thought of as an admissible e-variable, in the sense that no other e-variable uniformly dominates it, a majorising e-class is a complete class of e-variables, and the optimal e-class represents the minimal complete class. Notably, the idea of admissibility has already been explored in the e-variable literature. Indeed, Ramdas et al. (2022) considered a notion of admissible wealth processes, which in some contexts is closely related to the maximality for e-variables discussed here. However, their results do not suffice to characterise the optimal e-class in the setting that we study, as already pointed out by Clerico (2024). Admissibility has also been recently examined in a different context involving e-variables, in the problem of “merging” e-variables (Wang, 2024). Exploring the connections between those ideas and the concepts discussed in this work could be a promising direction for future research.
To establish our results, we employed a standard functional analytic approach. We began by proving the optimality of the dual e-class for a set $\mathcal{X}$ with finite cardinality, then extended it to the case of a compact $\mathcal{X}$ using a dense countable subset, and finally leveraged sigma-compactness for a general closed $\mathcal{X}$. Our proofs are self-contained and avoid relying on advanced tools from functional analysis or measure theory. We mention, however, that an alternative approach, leveraging bipolar-like theorems, might potentially yield similar conclusions, though it remains unclear at this stage whether it would achieve the desired results and simplify the analysis. Connections between e-variables and polar sets have already been exploited in Larsson et al. (2024), building upon recent technical advances in characterising the bipolar of subsets of probability measures (see, e.g., Langner and Svindland (2022) and the references therein). Exploring the potential application of this approach to our framework is an interesting direction for future work, which might also help addressing more general settings.
@@
Finally, open questions remain regarding the minimal conditions, on the hypothesis $\mathcal{H}$, necessary for the existence of the optimal e-class, and the extent to which such e-class can be explicitly characterised, when it does exist. A simpler, perhaps more approachable, problem is whether it is possible to extend our analysis to hypotheses defined by countably many scalar constraints, which would allow to determine the optimal e-classes to estimate the mean of random variables with a known bound on the moment generating function (e.g., sub-Gaussian random variables).

### v3

@@ 6 Domains with finite cardinality
Lemma 4 and Lemma 2 ensure that the optimal e-class exists. Hence, by Lemma 1, it is sufficient to show that $\mathcal{E}_{\mathcal{H}}^{\vee}$ coincides with the set of all maximal e-variables. By Lemma 9, all the elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal. Therefore, we are left to show that every maximal e-variable is in $\mathcal{E}_{\mathcal{H}}^{\vee}$. Fix a maximal e-variable $E$. We adopt the following proof strategy. First, we show that there is a fully supported $\tilde{P}\in\mathcal{H}$ such that $\langle\tilde{P},E\rangle=1$. Using that such $\tilde{P}$ is in the relative interior of $\mathcal{H}$, we establish that $\langle P,E\rangle=1$ for all $P\in\mathcal{H}$. Finally, we see that this property is enough to ensure that $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$.
@@
Since this is the maximum among a finite number of continuous functions, it is continuous. For $\varepsilon$ large enough, $\varphi_{x}(\varepsilon)>1$. Since $\varphi_{x}(0)<1$, there is $\tilde{\varepsilon}>0$ such that $\varphi_{x}(\tilde{\varepsilon})=1$. Let $\tilde{E}=E+\tilde{\varepsilon}\mathbf{1}_{x}$. For any $\hat{P}\in\mathcal{V}_{0}$ we have $\langle\hat{P},\tilde{E}\rangle=\langle\hat{P},E\rangle=1$, as $x\notin\operatorname*{Supp}(\hat{P})$. Moreover, for all $\hat{P}\in\mathcal{V}\setminus\mathcal{V}_{0}$, $\langle\hat{P},\tilde{E}\rangle\leq 1$, by definition of $\varphi_{x}$ and $\tilde{\varepsilon}$. Therefore, $\tilde{E}$ is an e-variable, since its expectation is at most one under all the vertices of $\mathcal{H}$. As $\tilde{\varepsilon}>0$, $\tilde{E}\succ E$, and so $E$ cannot be maximal, a contradiction. We conclude that $S=\mathcal{X}$.
@@ 7 Matching sets
By definition of compatible set, $\mathcal{H}_{S}$ is non-empty, and hence a hypothesis on $S$. We list here a few results outlining basic properties of the restrictions of finitely and properly constrained hypotheses that we will use in the next section. The proofs, rather short, can be found in Appendix A.
@@ 8 Optimality of the dual e-class
As all elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal (Lemma 9), by Lemma 1 it is enough to show that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. Fix $E\in\mathcal{E}_{\mathcal{H}}$. By Lemma 7 we can fix a minimal constraint $\Phi$ for $\mathcal{H}$. By Lemma 16, there is a finite matching set $S^{\star}\subseteq\mathcal{X}$ such that any closed set $S\subseteq\mathcal{X}$ containing $S^{\star}$ is a matching set inducing a compact $\Lambda_{\Phi|S}$. As $\mathcal{X}$ is closed, it is sigma-compact, and there is a sequence $(K_{i})_{i\geq 1}$ of compact matching sets, such that $S^{\star}\subseteq K_{i}\subseteq K_{i+1}$, for all $i\geq 1$, and $\mathcal{X}=\bigcup_{i=1}^{\infty}K_{i}$. By construction, $\mathcal{H}_{K_{i}}$ is properly constrained (Lemma 12) and $\Phi\big{|}_{K_{i}}$ a constraint for it (Lemma 11). From Proposition 5, for each $i\geq 1$ there is $\lambda_{i}\in\Lambda_{\Phi|K_{i}}$ such that $E_{\lambda_{i}}(x)\geq E(x)$ for all $x\in K_{i}$, where as usual $E_{\lambda}=1-\lambda\cdot\Phi$. For all $i\geq 1$, because $K_{i}\subseteq K_{i+1}$ we have $\Lambda_{\Phi|K_{i}}\supseteq\Lambda_{\Phi|K_{i+1}}$. In particular, for any $j\geq 1$, $(\lambda_{i})_{i\geq j}$ is contained in $\Lambda_{\Phi|K_{j}}$, which is compact by Lemma 16. So, there is a subsequence $(\lambda^{\prime}_{i})_{i\geq 1}$ of $(\lambda_{i})_{i\geq 1}$ converging to $\lambda^{\star}\in\bigcap_{j\geq 1}\Lambda_{\Phi|K_{j}}$. Let $(K^{\prime}_{i})_{i\geq 1}$ be the corresponding subsequence of $(K_{i})_{i\geq 1}$, obtained selecting the same indices. By Lemma 14, $\lambda^{\star}\in\Lambda_{\Phi}$, so $E_{\lambda^{\star}}\in\mathcal{E}_{\mathcal{H}}^{\vee}$. Fix any $x\in\mathcal{X}$. There is an index $i^{\star}\geq 1$ such that $x\in K_{i}^{\prime}$ for all $i\geq i^{\star}$. Thus, $E_{\lambda^{\prime}_{i}}(x)\geq E(x)$ for all $i\geq i^{\star}$, and $E_{\lambda^{\star}}(x)\geq E(x)$. As the choice of $x$ has been arbitrary, $E_{\lambda^{\star}}\succeq E$. Therefore, $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. ∎
@@ 9.2 Loosely constrained hypotheses
For any properly loosely constrained hypothesis the optimal e-class exists. This is made explicit in the next proposition (see Appendix C for the proof).
@@
The set =$\mathcal{H}=\{P\in\mathcal{P}_{\Phi}\,:\,\langle P,\Phi\rangle=0\}$, where $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ is a continuous function, is non-empty (and thus a hypothesis) if, and only if, $0\in\operatorname*{conv}\Phi(\mathcal{X})$ (Lemma 21 and Lemma 23). Conversely, when we allow for inequality constraints, as in the definition of loosely constrained hypotheses, this is not anymore a necessary condition. More explicitly, =$\mathcal{H}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\,:\,\langle P,\Phi^{\prime}\rangle=0\,,\,\langle P,\Phi^{\prime\prime}\rangle\geq 0\}$ can be non empty if $0\notin\operatorname*{conv}\Phi(\mathcal{X})$, (where $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ are continuous functions on $\mathcal{X}$ and Φ=($\Phi=(\Phi^{\prime},\Phi^{\prime\prime})$). However, here we only focus on the case $0\in\operatorname*{conv}\Phi(\mathcal{X})$, as required by Definition 10, and defer to future work the analysis of the setting of generic hypotheses defined via loose constraints. We also note that Lemma 18 cannot be trivially extended to the case of loosely constrained hypotheses, and it is not always possible to restrict the domain $\mathcal{X}$ of a loosely constrained hypothesis to make it properly loosely constrained. For an example where this is not possible, consider the case of a loosely constrained hypothesis on $\mathcal{X}=\mathbb{R}$, with tight constraint $\Phi^{\prime}:x\mapsto 1-x$ and slack constraint $\Phi^{\prime\prime}:x\mapsto x^{2}-1$.
@@ 10.1 Mean estimation for bounded random variables
$\mathcal{H}_{\mu}$ is a hypothesis on $\mathcal{X}$ (it is non-empty by Lemma 21). It is finitely constrained, with $\Phi_{\mu}:x\mapsto\mu-x$ a constraint (as $0\in\operatorname*{conv}\Phi_{\mu}(\mathcal{X})$ because $\mu\in\mathcal{C}$). Moreover, if $\mu\in\operatorname*{rel\,int}\mathcal{C}$, $\mathcal{H}_{\mu}$ is a properly constrained hypothesis on $\mathcal{X}$. Hence, for the sake of simplicity, we assume as known that $\mu^{\star}\in\operatorname*{rel\,int}\mathcal{C}$. Therefore, we are only interested in confidence sets $S_{t}\subseteq\operatorname*{rel\,int}\mathcal{C}$. We note, however, that the general case can be addressed by dealing with the values of $\mu$ on the relative boundary of $\mathcal{C}$ as discussed in Section 9.1. A valid confidence sequence for $\mu^{\star}\in\operatorname*{rel\,int}\mathcal{C}$ with confidence level $\delta\in(0,1)$ is $(S_{t})_{t\geq 1}$, where
@@
Here, $R_{t}(\mu)$ denotes the cumulative reward, up to round $t$, of a player who is playing a testing game as in Definition 2 on the null hypothesis $P^{\star}\in\mathcal{H}_{\mu}$, observing the sequence $(X_{i})_{i\geq 1}$. From the fact that $\{\mu^{\star}=S_{t}\,,\,\forall t\geq 1\}\iff\{R_{t}(\mu^{\star})\leq\log(1/\delta)\,,\,\forall t\geq 1\}$, which happens with probability at least $1-\delta$ by Proposition 1, it follows that $(S_{t})_{t\geq 1}$ is a confidence sequence.
@@ 10.2 Mean estimation for heavy-tailed distributions
where we note that the fact that $\langle P,X^{2}\rangle\leq 1$ automatically implies that $\langle P,X\rangle$ is well defined. Clearly, $\mathcal{H}_{1}=\{\delta_{1}\}$ and $\mathcal{H}_{-1}=\{\delta_{-1}\}$. It is easy to check that, for $\mu\in(-1,1)$, $\mathcal{H}_{\mu}$ is a loosely properly constrained hypothesis on $\mathbb{R}$, where $\Phi^{\prime}_{\mu}:x\mapsto\mu-x$ is a tight constraint and $\Phi^{\prime\prime}:x\mapsto 1-x^{2}$ is a slack constraint. Denoting as $R_{t}(\mu)$ the cumulative reward (up to round $t$) of a player betting in a testing game on $\mathcal{H}_{\mu}$ and observing the sequence $(X_{i})_{i\geq 1}$, reasoning as in Section 10.1, we deduce that
@@
By Proposition 6, for $\mu\in(-1,1)$ the optimal e-class is
@@
The problem of finding confidence sequences for heavy-tailed random variables via betting strategies has been addressed in the literature. Agrawal et al. (2021) obtained anytime-valid lower bounds for real random variables with bounded $(1+\varepsilon)$-th (non-centred) moment, via an approach that implicitly uses e-variables. Their Lemma 17 is in fact a regret upper bound for the cumulative reward of a player in a testing game as in Definition 2. Notably, our results show that their analysis considered games restricted to the optimal e-class. On the other hand, Wang and Ramdas (2023) leveraged ideas from Catoni (2012) to set up testing games in order to find confidence sequences for the mean of real random variables with bounded $(1+\varepsilon)$-th central moment. The theory that we have developed in this work shows that their choice of restriction for the e-variables does not correspond to the optimal e-class.
@@ 11 Conclusion
First, the key concepts of maximal e-variable and majorising and optimal e-class that we discuss in this work are strongly connected to the classical statistical notions of admissibility and completeness (see, e.g., Section 1.8 in Lehmann and Romano, 2022). Indeed, with this terminology, a maximal e-variable can be essentially thought of as an admissible e-variable, in the sense that no other e-variable uniformly dominates it, a majorising e-class is a complete class of e-variables, and the optimal e-class represents the minimal complete class. Notably, the idea of admissibility has already been explored in the e-variable literature. Indeed, Ramdas et al. (2022) considered a notion of admissible wealth processes, which in some contexts is closely related to the maximality for e-variables discussed here. However, their results do not suffice to characterise the optimal e-class in the setting that we study, as already pointed out by Clerico (2024). Indeed, it would be possible to use their results if we were considering hypotheses where a common reference measure (namely a measure for which every element in $P$ admits a density) exists, which is not the case in the setting we study. Admissibility has also been recently examined in a different context involving e-variables, when “merging” e-variables (Wang, 2024). Exploring the connections between those ideas and the concepts discussed in this work could be a promising direction for future research.
To establish our results, we employed a standard functional analytic approach. We began by proving the optimality of the dual e-class for a set $\mathcal{X}$ with finite cardinality, then extended it to the case of a compact $\mathcal{X}$ using a dense countable subset, and finally leveraged sigma-compactness for a general closed $\mathcal{X}$. Our proofs are self-contained and avoid relying on advanced tools from functional analysis or measure theory. Building on the formulation and findings presented in the first pre-print of this work, Larsson et al. (2025) have recently re-derived our characterization of the optimal e-class for constrained hypotheses and have extended it to a significantly more general framework. To achieve this broader generality, their approach relies on sophisticated topological and functional-analytic techniques to derive duality results in functional spaces. Nonetheless, we believe that in scenarios where the problem can be reduced to finite-dimensional analysis, our more straightforward and elementary method remains valuable and more accessible.
As highlighted in the introduction, our analysis is restricted to the characterisation of optimality within the testing-by-betting framework with e-values, as defined in Definition 2. More generally, however, sequential testing with e-values may proceed differently. At each round $t$ one might select an e-variable $E(x_{1},\dots,x_{t})$, with respect to a hypothesis concerning the entire sequence of observations up to time $t$. In such settings, it is natural to define the hypothesis $\mathcal{H}$ on the space of stochastic processes taking values in $\mathcal{X}$, rather than on $\mathcal{X}$ itself. This broader perspective is, for instance, essential when testing whether a sequence of observations is i.i.d. These kinds of sequential tests typically rely on e-processes, namely non-negative stochastic processes whose expected value remains bounded by one, under the null hypothesis, for any stopping time (Ramdas and Wang, 2024; Ramdas et al., 2022). In contrast, the present work focuses exclusively on single-round e-variables, where the hypothesis is defined on on $\mathcal{P}_{\mathcal{X}}$, the space of distributions over a single observation. This framework can be appropriate when one assumes that the observed data arise from i.i.d. draws from a $P\in\mathcal{P}_{\mathcal{X}}$, and is interested solely in verifying whether this $P$ is within $\mathcal{H}$. In this sense, tests that might reject the hypothesis on the basis that the data appear non-i.i.d. are not considered, as the i.i.d. assumption is given for granted (one might want to formalise this by asking for test that has power under any alternative stating that the data are i.i.d.). The more general problem of characterising the optimal class of e-processes for suitably constrained hypotheses (for instance under i.i.d. assumptions, or assuming that constraints hold conditionally) is left open for future investigation.
@@
Finally, open questions also remain regarding the minimal conditions, on the hypothesis $\mathcal{H}$, necessary for the existence of the optimal e-class, and the extent to which such e-class can be explicitly characterised, when it does exist.